INAF LBC+MODS+LUCI queue observing, 2012 October 24-25

Observer: M. Faccini, E. Sani, K. Boutsia
Telescope Operator: G. Bechetti
Support Astronomer: M. Edwars
Instrument: LUCI, LBC

Summary

At the beginning of the twilight we have clear sky with wind gusts at 15-18m/s
At the beginning of the night the seeing is 1.7 and the wind is lowering
During the spectroscopic acquisition of SN2009ip we have problems with collimation. The target is too low (<30elev) and the seeing is high. We observe nevertheless because the PI asked the data under any conditions.
Before starting the second filter (HKspec), we take a through-slit image and the target seems to still be in the slit. We go on even though collimation gets worse. Guiding also fails and we have tracking
At the end of this project, the collimation is so bad that we have to call D.Miller. Trying to re-collimate LUCI, the Cresci target goes out of specs and we switch to LBC.
LBC was more stable tonight. Even though we rolled back to build-2012F, the elongation problem was not completely solved and we had to switch off guiding again at some point.
During twilight we tried to obtain 2 standards, but at the second one the background is too high. Also twilight flats are somewhat hight but they should be usable.

Data Summary

Project_RUN Objects Filter ExpTime Notes
LaFranca_6 3C403 zJspec, 200_H+K 12m Test to check the slit alignment and the saturation of the reference star
Turatto_7A SN2009ip zJspec, 200_H+K 1hr Check of the slit alignment. No collimation. flat+arc+Tell - COMPLETE
Turatto_7A SN2009ip HKspec, 200_H+K 45m No collimation. No guiding, flat+arc+Tell - COMPLETE
Giallongo_A MS0015 g-sloan, r-sloan 30min out of spec
Giallongo_B MAC0717 g-sloan, r-sloan 2hr after allseeing we have median seeing of 1.2" for 1hr15min of blue / red: 1.0" - COMPLETE
Standard SA98 g-sloan, r-sloan   airmass 1.2
Standard SA101 g-sloan, r-sloan   airmass 1.4 - background too high, not sure it is usable

Details

00:52 We start with flats calibrations for the J, H, Ks filters
  • execuLUCIScript.sh flatKs.txt luci.0005-0034
  • execuLUCIScript.sh DoFlats_sunset.txt luci.0035-0119

01:49 We slew to La Franca A15
  • execuLUCIScript.sh OB_LaFranca_A15_acq.txt luci.0120-0126

01:54 execuLUCIScript.sh OB_LaFranca_A15_spe.txt luci.0127-0128
  • this is a test to verify whether the reference/telluric star saturate. The star does not saturate.
  • Perhaps the centering of the slit is not perfect and require to be refined through a small field rotation.
  • A standard star is already within the slit.

02:15 We slew to Turatto 7B
  • we start with the imaging to verify the offsets.
  • execuLUCIScript.sh SN2009ip_oct2012_H.img luci.0129-0140
  • We abort the imaging because we can not see the target on the acquisition image. This is problably due to a bad seeing of 2.5

2:25 We slew to the spectroscopic program Turatto 7A
  • execuLUCIScript.sh SN2009ip_ls.acq luci.0141-0144

02:36 execuLUCIScript.sh SN2009ip_zJ.spec luci.0145-0156
  • The collimation is not good

3:43 We try to check the guiding luci.0157-0162
  • Even if ther is no collimation, we verify that the target is still in the slit.

3:49 execuLUCIScript.sh SN2009ip_HK.spec luci.0163-0171
  • We loose the guide star and have to restart the script. This happens twice.

04:45 We perform the imaing in the Ks-band.
  • execuLUCIScript.sh SN2009ip_oct20012_Ks.img
  • We are at 25.3deg in Elevation, there is no way to guide. NO Ks imaging
  • The J- and H- bands imaging was performed during the acquisitions of the target.

04:51 We slew to the Telluric Hip117452.

05:05 We can not point any target. We have Dug online

05:44 Manual collimation.
  • execuLUCIScript.s tell_HK.acq luci.0171-0175

05:52 execuLUCIScript.sh tell_HK.spec luci.0176-0181
  • execuLUCIScript.sh tell_zJ.spec luci.0182-0187
  • We need to change the telluric because of saturation

6:01 We slew to Hipp104320 with the Turratto specs
  • execuLUCIScript.sh tell_back.acq luci.0188-0190
  • execuLUCIScript.sh tell_zJ.spec luci.0191-0196

06:18 execuLUCIScript.sh tell_HK.spec luci.0197-0202
  • The spectrum appear too faint, we double the DIT

06:22 execuLUCIScript.sh tell_HK.spec luci.0203-luci.0208

6:27 We repeat Hipp104320 with the Cresci specs
  • execuLUCIScript.sh tell_acq_campoA.txt luci.0209-0212
  • execuLUCIScript.sh tell_int_campoA_H.txt ==luci.0213-0220

06:52 We abort Cresci A program because we are no able to collimate.
  • We ask the TO to slew to a bright star near Giallongo 27A and collimate to measure the seeing with the DIM.
  • Nonetheless we have problems with the LBC software.

07:11 We switch to LBC

07:35 Slew to Pointing Star in Giallongo 27A for focus and lbcrangebal
  • dofpia, /first, X2
  • dofpia, /X2, after fiew iteration converge
  • lbcrangebal B=075113 R=075107

07:57 We start MS0015.9+16_GR_d.ob
  • We are out of the specs (close to the moon, seeing 1.8), but this is the only target available for now.

08:43 We slew to Giallongo 27B, run of MAC0717+37_GR_POINTING.ob
  • dofpia
  • lbcrangebal B=085243 R=085241

9:02 Start MAC0717+37_GR_d.ob

9:10 End of MAC0717+37_GR_d.ob
  • dofpia

9:40 Start MAC0717+37_GR_e.ob

10:15 Fringing pattern appear in r08101003

10:31 End of MAC0717+37_GR_e.ob
  • dofpia

10:38 Start of MAC0717+37_GR_f.ob
  • We exclude the guiding

11:20 End of MAC0717+37_GR_f.ob
  • dofpia

11:23 Re-start of MAC0717+37_GR_f.ob
  • We exclude the guiding

11:47 Stop of MAC0717+37_GR_f.ob due to elongation in both B- and R- channel
  • dofpia

11:53 Re-start MAC0717+37_GR_f.ob
  • We exclude the guiding

12:02 Abort due to transit (bad tracking)
  • dofpia

12:06 Re-start MAC0717+37_GR_f.ob
  • We exclude the guiding

12:19 Stop of MAC0717+37_GR_f.ob due to elongation in both B- and R- channel
  • dofpia

12:25 Re-start MAC0717+37_GR_d.ob
  • dofpia

12:36 Abort due to increasing background

12:37 We slew to SA98 (airmass 1.22)

12:51 We slew to Sa101 (airmass 1.40)
  • The background is too high

12:55 We start sky flats. The couts are a bit high (20000-40000 cts)

01:10 We close the dome.

01:25 LUCI CALIB
  • execuLUCIScript.sh OB_LaFranca_CAL_flat.txt luci.0231-0249
  • execuLUCIScript.sh HKspec.flat luci.0251-0260 The frame luci.0250 is not in the /Repository/20121025/
  • execuLUCIScript.sh HKspecLong.flat luci.0261-0270
  • execuLUCIScript.sh HKspec.arc luci.0270-0274 The frame luci.0272 is not in the /Repository/20121025/
  • execuLUCIScript.sh HKspecLong.arc luci.0275-0286
  • execuLUCIScript.sh zJspec.arc luci.0287-0290
  • execuLUCIScript.sh zJspecLong.arc luci.0291-0302
  • execuLUCIScript.sH OB_LaFranca_CAL_lamp.txt luci.0303-0314
  • execuLUCIScript.sH dark_int.txt luci.0315-0333
  • execuLUCIScript.sH dark_nor.txt ==luci.0334-

END OF NIGHT - END OF RUN

-- RobertoSpeziali - 25 Oct 2012
Topic revision: r9 - 28 Oct 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback